Vicky is 4 times as old as Ella Vicky is 15 years older then Ella how old are they both

Answer by ikleyn(52794)

.

V = 4E       (Vicky is 4 times as old as Ella)

V - E = 15   (Vicky is 15 years older than Ella)


Hence

    4E - E = 15

    3E     = 15

     E     = 15/3 = 5.


ANSWER.  Ella is 5 years old;  Vicky is 4*5 = 20 years old.    ANSWER

Solved.